OpenStudy (anonymous):

Sigma Notation: Need Help!! Problem 1: Evaluate the summation of 3 times negative 2 to the n minus 1 power, from n equals 1 to 5. Choices: A.-93 B.-33 C.33 D.93

OpenStudy (anonymous):

I don't know how to do this and would appreciate help and an explanation.

OpenStudy (anonymous):

\[\sum_{n=1}^{5} 3(-2) ^ n-1\]

OpenStudy (dumbcow):

there may be a faster way but as far as i can see, just plug in the "n" values and add up the results

OpenStudy (anonymous):

I don't know how :/

OpenStudy (dumbcow):

the 3 can go on outside since its just a multiplier \[\rightarrow 3\sum_{n=1}^{5}(-2)^{n-1}\]

OpenStudy (anonymous):

The way I wrote it is how it is set up

OpenStudy (dumbcow):

well what is (-2)^0? (-2)^1 and so on

OpenStudy (dumbcow):

oh sorry so the "-1" is not part of the exponent?

OpenStudy (aum):

\[ \sum_{n=1}^{5} 3(-2) ^ n - 1 = \\ \sum_{n=1}^{5} 3(-2) ^ n - \sum_{n=1}^{5}1 = \\ 3\sum_{n=1}^{5} (-2) ^ n - 5 \\ 3(-2 + 4 - 8 + 16 - 32) - 5 = ? \]

OpenStudy (anonymous):

not that I am aware of I wrote it as it is written. I don't know the steps or how to solve it..

OpenStudy (dumbcow):

hmm that give -71 which is not an answer choice i think its meant to be as I wrote it

OpenStudy (anonymous):

I don't know. its on my work which I have like 4 probs like this involving sigma notation and that's exactly how it is written.

OpenStudy (aum):

\[ \sum_{n=1}^{5} 3(-2) ^ {n - 1} = 3\sum_{n=1}^{5} (-2) ^{n-1} = \\ 3(1 - 2 + 4 - 8 + 16) = 3 = 33 \]

OpenStudy (anonymous):

There is also this: Evaluate the summation of 2 n plus 5, from n equals 1 to 12. Choices: A.29 B.36 C.216 D.432 @aum

OpenStudy (aum):

You have to clarify if 2 is raised to an exponent or 2 is multiplied by n. Is 5 in the exponent?

OpenStudy (aum):

Use ^ for exponents. Use parethesis to group items if necessary.

OpenStudy (anonymous):

I have two more after this one if you could help. The one I just posted looks like this: \[\sum_{n=1}^{12} 2n+5\]

OpenStudy (aum):

\[ \sum_{n=1}^{12}2n+5 = \sum_{n=1}^{12}2n + \sum_{n=1}^{12}5 = \\ 2\sum_{n=1}^{12} n + (5 * 12) = 2\frac{12*13}{2} + 60 = ? \]

OpenStudy (anonymous):

How do I do math to solve top one. I know 5*12=60 and 12*13/2 = 78 then when you add 60 it is 138

OpenStudy (aum):

The 2 on the top and bottom will cancel. Just (12 * 13) + 60

OpenStudy (anonymous):

so 216? is that the answer? and can you help and explain with two more?

OpenStudy (aum):

For the above problem I have used the formula for adding: 1 + 2 + 3 + 4 + ..... + n = n(n+1) / 2 Here n = 12. So 12 * 13 / 2 is the sum To that sum add 60 to get 216. Yes, 216.

OpenStudy (aum):

go ahead.

OpenStudy (anonymous):

\[\sum_{n=2}^{10}25(0.3)^n+1\] n+1 is small

OpenStudy (aum):

What do you mean n+1 is small. Are both n and 1 in the exponent?

OpenStudy (anonymous):

yes exactly like their the power of or to n+1

OpenStudy (aum):

Then you should write like this: \[\large \sum_{n=2}^{10}25(0.3)^{n+1} \]

OpenStudy (anonymous):

yes exactly how it looks I just didn't know how to make both small. Only the n is what I knew how to do.

OpenStudy (aum):

Right click on the formula I wrote and click on "Show Math as" and then "Tex commands". It will show the exact latex code that anyone uses to write their stuff. Just the n+1 should be enclosed in {n+1} to have both as exponents.

OpenStudy (anonymous):

oh okay I gotcha so how would we set this up to solve this one. After this one theres one more then you can go, I appreciate the help!!

OpenStudy (aum):

\[ \large \sum_{n=2}^{10}25(0.3)^{n+1} = 25\sum_{n=2}^{10}(0.3)^{n+1} = \\ \large 25[0.3^3 + 0.3^4 + 0.3^5 + .... +0.3^{11}] = 25 * [~~??~~] \] Use the formula for the sum of a geometric sequence to add up \(\large [0.3^3 + 0.3^4 + 0.3^5 + .... +0.3^{11}]\)

OpenStudy (anonymous):

-Aum can you please help me with sample size

OpenStudy (anonymous):

is it 3.3? @aum

OpenStudy (aum):

What are your answer choices?

OpenStudy (anonymous):

A.0.005 B.0.321 C.0.964 D.10.714

OpenStudy (anonymous):

Sample size of +- 7%

OpenStudy (anonymous):

I got 3.0375 I don't get it/: what did I do wrong? or what am I not doing?

OpenStudy (aum):

Use the formula for the sum of a geometric sequence to add up \(\large [0.3^3 + 0.3^4 + 0.3^5 + .... +0.3^{11}]\) The formula for adding n terms of a geometric sequence, a + ar + ar^2 + ....+ar^(n-1) is: a * (1 - r^n) / (1 - r). We need to add 0.3^3 + 0.3^4 + 0.3^5 + ..... + 0.3^(11) Here the first term a = 0.3^3 common ratio r = 0.3 number of terms when n goes from 2 to 10 is 9. Plug the numbers into the geometric sum formula. Then multiply the sum by 25 to get the answer.

OpenStudy (anonymous):

Bruh!

OpenStudy (anonymous):

sorry im trying to add them up but its not making sense, when I get to multiply 25 it hikes up to 29

OpenStudy (aum):

\[ \large 25[0.3^3 + 0.3^4 + 0.3^5 + .... +0.3^{11}] = 25 * 0.3^3\frac{1 - 0.3^9}{1-0.3} = \\ \large 25 * 0.027 *\frac{0.99998}{0.7} = ? \]

OpenStudy (anonymous):

The last problem I ll need help with will look like this if you wanna look at it while I figure this one out. \[\sum_{n=3}^{12} -5n-1\] And choices are: A.-410 B.-385 C.410 D.385

OpenStudy (anonymous):

is it C.0.964?

OpenStudy (aum):

\[ \large \sum_{n=3}^{12} -5n-1 = -5\sum_{n=3}^{12} n -\sum_{n=3}^{12} 1 = \\ \large -5 * \frac{(3+12)}{2} * 10 - 10 = ? \]

OpenStudy (aum):

Yes, C 0.964 for the previous problem.

OpenStudy (anonymous):

-385? right?:D

OpenStudy (aum):

yes.

OpenStudy (anonymous):

WOO thank you @aum your a life saver!

OpenStudy (aum):

You are welcome.